OpenStudy (anonymous):

Which of the following ordered pairs is not a solution to the inequality y greater than or equal to -4x - 2? (-1, 0) (0, 0) (4, 4) (1, 3)

OpenStudy (anonymous):

Try them. Remember (x,y) (-1, 0) = > -4(-1)-2 = 4-2 = 2 0 < 2, this is NOT a solution (0,0) => -4*0 -2 = -2 0 > -2 This is a solution! try the others